19 gm mixture of Na2CO3 and NaHCO3 on heating gives 2.2 gm CO2 gas. Find % NaHCO3 by mass in mixture.​..

Grade:12th pass

1 Answers

Pawan Prajapati
3 years ago
Answer:

2NaHCO3

Δ

Na2CO3

+H

2

O+CO

2

1.12 L CO

2

at STP corresponds to

22.4L/mol

1.12L

=0.05 moles.

0.05 moles of CO2

will come from 2×0.05=0.1 mole of NaHCO3

.

The molar mass of NaHCO3

is 84 g/mol. Mass of 0.1 moles of NaHCO3

=84×0.1=8.4g

Mass of Na2CO3

=19−8.4=10.6g

Thanks